# Supplementary Materials

### v1

@@ 4.4.2 $G$ is Unknown: Symmetry Discovery Framework
For invariance tasks, we use only the $\phi_{n}$ network and omit $\phi_{s}$. When the group acts on the input domain via a product action, we adopt the function decomposition from (9). This decomposition is applied both when the group is explicitly known—where we incorporate it into (10)—and when the group is unknown, in which case it is used in (11).
Method Orthogonal Lorentz Arbitrary
$G$-Ortho-Nets $1.12\times 10^{-6}$ $8.72\times 10^{-7}$ $6.64\times 10^{-7}$
EMLP $6.64\times 10^{-2}$ $4.66\times 10^{-2}$ N/A
Simple MLP $7.83\times 10^{-5}$ $1.83\times 10^{-5}$ $2.41\times 10^{-5}$
Table 1: MSE on validation dataset using G-Ortho Nets is evaluated for Orthogonal $O(n)$, Lorentz $O(1,3)$, and Arbitrary Group for invariance task. To ensure a rigorous comparative analysis, all models are designed with an identical parameter count and trained for the same number of epochs, ensuring a fair and consistent evaluation. Note: EMLP failed to converge for the Arbitrary Group.
Method Moment of Inertia Top Tagging
Val Loss Accuracy AUROC
$G$-Ortho-Nets $1.16\times 10^{-5}$ $0.915$ $0.969$
EMLP $1.36\times 10^{-3}$ N/A N/A
LieGNN N/A $0.938$ $0.984$
Table 2: Performance Comparison on Moment of Inertia Prediction and Top Tagging Classification Tasks: The table shows that while G-Ortho Nets achieves slightly lower accuracy than LieGNN, it unifies symmetry discovery and classification in a single network. In contrast, LieGNN employs a two-stage approach using LieGAN for symmetry discovery followed by a separate GNN for classification.
Task $G$-Ortho-Nets LieGAN
$G$-Invariant Polynomial Regression $9.99$ $9.99$
Top Tagging $9.99$ $9.99$
Equivariant Polynomial Regression $9.99$ $9.99$
Moment of Inertia $9.99$ $9.99$
Table 3: Comparison of cosine similarity between $G$-Ortho-Nets and LieGAN across multiple tasks.
(a) GTGT
(b) LieGANLieGAN
(c) Ours
(d) GTGT
(e) LieGANLieGAN
(f) Ours
Figure 3: Comparison of ground truth and learned matrices. (Top) Top quark tagging: the ground truth matrix $A=\operatorname{diag}(1,-1,-1,-1)$ is compared with matrices learned by LieGAN and $G$-Ortho-Nets . (Bottom) Invariant polynomial regression with an arbitrary matrix AA: the ground truth is compared with matrices recovered by LieGAN and $G$-Ortho-Nets. Both methods accurately recover the underlying matrix structure.
5 Experiments and Results
We evaluate our proposed framework on both $G$-invariant and $G$-equivariant tasks, where $G$ is an orthogonal group as defined in Eq. (6). To assess its performance, we consider three distinct types of groups: the standard orthogonal group, the Lorentz group, and a group that preserves a random quadratic form $x^{T}Ax$, where $A$ is a randomly generated symmetric matrix. Across all experiments, we compare the following model architectures:
  • G-Ortho-Nets: Our proposed method, which incorporates the product of a norm-invariant and a scale-invariant function.
  • EMLP: EMLP architecture with G-equivariant linear and bilinear layers as discussed in (Finzi et al., 2021).
  • LieGAN: Generative Adversarial Symmetry Discovery proposed by (Yang et al., 2023)
  • LieGNN: LieGAN integrated with a graph neural network.
  • Simple MLP: A standard feed-forward multilayer perceptron.
The experimental setup is detailed in the subsequent subsections.
5.1 $G$-Invariant Polynomial Regression
(a) GTGT
(b) LieGANLieGAN
(c) Ours
Figure 4: Comparison of the ground truth matrix AA (first plot) with matrices learned by LieGAN (second plot) and $G$-Ortho-Nets (third plot) for the $G$-Equivariant Polynomial Regression task.
We conduct a series of experiments to evaluate our method’s ability to learn invariant polynomial functions of the form:
$f(x)=\sin^{2}(x^{T}Ax)+\sin(x^{T}Ax)$ (13)
where the invariance structure is determined by the matrix $A$. We examine three distinct settings, each characterized by a different choice of $A$, to demonstrate the versatility and robustness of our approach.
Experimental Setting. We evaluate our method on three group-invariant polynomial regression tasks. For the orthogonal group, we set $A=I_{4}$ to construct an $O(4)$-invariant task. For the Lorentz group, we employ $A=\text{diag}(1,-1,-1,-1)$, corresponding to the Minkowski metric. For the arbitrary group case, we use a randomly chosen matrix $A$ to demonstrate the methods ability to generalize beyond well-studied groups in the literature. All tasks use the same target polynomial function.
Results. Table LABEL:tab:1 shows that $G$-Ortho Nets achieves lower MSE on the validation dataset compared to EMLP and Simple MLP baselines. Figures LABEL:fig:orthogonal_matrix6, and LABEL:fig:arbitrary_matrix illustrate that our method successfully recovers the underlying invariant structure, with learned matrices closely matching the ground truth across all three group settings. These results demonstrate the effectiveness of our method in both structure discovery and function approximation tasks.
5.2 Lorentz Invariance: Top Tagging
Experimental Setting. We evaluate our method on Lorentz-invariant top tagging task using the Top Quark Jet Dataset (Kasieczka et al., 2019). The task requires classifying jets originating from top quarks versus those from lighter quarks while preserving Lorentz symmetry. The underlying Lorentz metric $A=\text{diag}(1,-1,-1,-1)$ governs the symmetry structure, which our method discovers jointly with the classification task.
Results. $G$-Ortho Nets achieves comparable performance to LieGAN on the top tagging task, as shown in Table 2, with similar accuracy and AUROC scores. Additionally, Figure 6 demonstrates that our method successfully recovers the underlying Lorentz metric structure $A=\text{diag}(1,-1,-1,-1)$.
5.3 Equivariant Polynomial Regression
Experimental Setting. We evaluate our method on $G$-equivariant polynomial regression task with arbitrary group structure defined by matrix $A$. The target polynomial is
$f(x)=9(xx^{T}A)^{2}+2(xx^{T}A),$ (14)
which is equivariant with respect to $G$, since $f\mathopen{}\left({Ux}\,\negthickspace\right)\mathclose{}=Uf(x)U^{-1},\;\forall U\in G$.
Results. Figure LABEL:fig:equivariant_matrix compares the ground truth matrix $A$ with reconstructions from LieGAN and $G$-Ortho Nets. Table LABEL:tab:1 shows that both methods accurately recover the underlying group structure based on cosine similarity. However, $G$-Ortho Nets integrates symmetry discovery with equivariant function learning in a single framework, whereas LieGAN performs only structure discovery. This enables our method to simultaneously recover the group structure and learn the target polynomial function.
5.4 $O(3)$-Equivariance: Moment of Inertia
Experimental Setting. We evaluate our method on predicting the moment of inertia $\mathcal{I}$ from particle masses and positions $p=\{(x_{i},m_{i})\}_{i\in[M]}$. The moment of inertia is computed as
$\mathcal{I}=\sum_{i}m_{i}\left(x_{i}^{T}x_{i}\,I-x_{i}x_{i}^{T}\right),$ (15)
which is $O(3)$-equivariant. We compare our method against EMLP (Finzi et al., 2021), which requires known group structure, and the symmetry discovery method from (Yang et al., 2023).
Results. Figure LABEL:fig:moi_matrix compares the ground truth matrix $A=I_{3}$ (normalized) with reconstructions from LieGAN and $G$-Ortho-Nets. Both methods successfully recover the $O(3)$-equivariant structure. Table 2 shows that $G$-Ortho-Nets achieves lower MSE than EMLP on the prediction task. Table LABEL:tab:1 reports cosine similarity of 0.99 for both LieGAN and $G$-Ortho-Nets, indicating accurate structure recovery.
@@ 5.5 Limitations
Our method is currently restricted to generalized orthogonal groups and supports only standard group actions and diagonal (product) group actions. In future work, we aim to overcome these limitations by extending our framework to accommodate a broader range of symmetry groups and group actions.
@@ 6 Conclusion
In this work, we introduced a framework for learning group-equivariant functions using quadratic forms, with a focus on orthogonal groups. We showed that such functions can be expressed in a canonical form, incorporating both scale-invariant and norm-invariant components. Additionally, we extended our approach to a more general setting where the function acts on tuples of input vectors under a diagonal (or product) group action. In this extension, the equivariant function is decomposed into an angular component derived from the normalized first vector and a scale-invariant component based on the full Gram matrix of the tuple. This decomposition effectively captures inter-dependencies between multiple inputs while maintaining group symmetry.
Through extensive experiments, we demonstrated the ability of our method to discover underlying symmetries and efficiently learn equivariant functions across various tasks. In future work, we aim to extend our approach to broader symmetry groups beyond the general orthogonal group and explore a wider range of group actions.
Appendix A Appendix
(a) GTGT
(b) LieGANLieGAN
(c) $G$-Ortho-Nets
(d) GTGT
(e) LieGANLieGAN
(f) $G$-Ortho-Nets
Figure 5: Comparison of ground truth and learned matrices. Top $O(4)$-Invariant Polynomial Regression: The ground truth matrix $A=I_{4}$ is compared with matrices learned by LieGAN and $G$-Ortho-Nets. Both methods accurately recover AA (matching up to the first decimal) while modeling the polynomial $\sin^{2}(x^{T}Ax)+\sin(x^{T}Ax)$. Bottom $O(3)$-Equivariant Moment of Inertia: The ground truth matrix $A=I_{3}$ is compared with matrices learned by LieGAN and $G$-Ortho-Nets.
(a) GTGT
(b) LieGANLieGAN
(c) $G$-Ortho-Nets
Figure 6: Lorentz-Invariant Polynomial Regression: Ground truth $A=\operatorname{diag}(1,-1,-1,-1)$ (first plot) is compared with matrices learned by LieGAN (second plot) and $G$-Ortho-Nets (third plot). Both methods approximately recover AA and model the polynomial $\sin^{2}(x^{T}Ax)+\sin(x^{T}Ax)$.
You may include other additional sections here.
Appendix B Additional Results and Complete Proofs

### v2

@@ 4.4.2 $G$ is Unknown: Symmetry Discovery Framework
For invariance tasks, we use only the $\phi_{n}$ network and omit $\phi_{s}$. When the group acts on the input domain via a product action, we adopt the function decomposition from (8). This decomposition is applied both when the group is explicitly known, where we incorporate it into (9), and when the group is unknown, in which case it is used in (10).
5 Discussion
5.1 Existing Canonical Forms
We would like to highlight that, for $O(n)$, there exist various standard forms for equivariant functions, depending on the output tensor types and the group action. These forms can be adapted for generic orthogonal groups. However, they change depending on the tensor type and the group action, unlike our method, which provides a common form irrespective of the tensor type and the group action.
5.2 Uniqueness and Householder Reflection
The proposed canonical decomposition of a $G$-equivariant function is unique up to a group transformation. Furthermore, under such a transformation, the output of the scale-invariant component of the function is equivalent to a Householder reflection transformation, $R$. This can be expressed as:
$\Phi_{s}\left(\frac{x}{\|x\|_{A}}\right)=\tau\left(I-\frac{2ww^{T}A}{w^{T}Aw}\right)\tau^{-1}$ (12)
where $\tau$ is an appropriate transformation that maps the group $G$ to the standard indefinite orthogonal group $O(p,q)$.
A key property of a Householder reflection is that it is its own inverse (i.e., $R=R^{-1}$). This insight leads to a significant simplification of our proposed form, particularly when the group action on the function’s output is a conjugation:
$f(g\cdot x)=g\cdot f(x)=gf(x)g^{-1}$ (13)
This simplification is formalized in the following proposition.
Proposition 5.1.
Let $f$ be a $G$-equivariant function where the group action on the output is a conjugation, as defined in eq. (13). Then, the function $f(x)$ can be expressed in the symmetric form:
$f(x)=\Phi_{s}\left(\frac{x}{\|x\|_{A}}\right)\Phi_{n}\left(\|x\|_{A}\right)\Phi_{s}\left(\frac{x}{\|x\|_{A}}\right)$ (14)
6 Experiments
We evaluate $G$-Ortho-Nets, our proposed method for learning $G$-equivariant functions through decomposition into norm-invariant and scale-invariant components as given in eq. (7). The quadratic form $A$ is learned jointly with the decomposed functions, enabling simultaneous discovery of the underlying symmetry group $O(p,q)$ which is defined in eq. (5).
6.1 Baseline Methods
Table 1: Suitability of related symmetry discovery methods for our setting. Only Augerino and LieGG provide both a usable prediction model and applicability to continuous equivariance tasks, motivating our choice of baselines. Our method further distinguishes itself by offering an interpretable model that trains in a single stage. ✓= supported, ×= not supported/limited.
Method Prediction Continuous Interpretable Limitations/Notes
Model Equivariance Prediction Model
G-Ortho-Nets (Ours) Interpretable prediction model, single-stage
Augerino × Sample averaging required at inference
LieGG × Post-training analysis needed
LieGAN × × No prediction model
LIG × × Requires synchronized output flows, two stage
SDBA × × Limited to single generator, two stage
BNN × Appropriate for discrete groups
Selected as baselines (only methods suitable for continuous equivariance with prediction models)
We compare $G$-Ortho-Nets against two recent approaches for symmetry discovery:
LieGG
(Moskalev et al. (2022)) retrieves infinitesimal generators of symmetries by solving a matrix nullspace equation derived from Lie group theory. Using singular value decomposition of the polarization matrix, it extracts the Lie algebra basis and quantifies symmetry bias and symmetry variance, measuring how closely learned generators approximate true underlying symmetries.
Augerino
(Benton et al. (2020)) learns invariances and equivariances by parameterizing a distribution over data augmentations. The model jointly optimizes network and augmentation parameters to discover symmetries directly from training data. Unlike LieGG’s post-hoc extraction, Augerino induces soft equivariance during training.
The rationale for selecting these two methods as baselines is summarized in Table 1.
6.2 Experimental Tasks
We evaluate on three tasks spanning synthetic regression, physical simulation, and high-energy physics applications. Table 2 summarizes the tasks and their corresponding symmetry groups.
Table 2: Experimental tasks and symmetry groups.
Task Group
Synthetic regression $O(2,2)$
Moment of inertia $O(3)$
Top quark tagging $O(1,3)$
6.2.1 Task 1: Synthetic Regression
We construct a target function exhibiting equivariance under the indefinite orthogonal group $O(2,2)$:
$f(x)=9\left(xx^{\top}Axx^{\top}A\right)+2\left(xx^{\top}A\right),$ (15)
where $A\in\mathbb{R}^{4\times 4}$ is symmetric and indefinite with signature (2,2)(2,2)—two positive and two negative eigenvalues. The function satisfies $f(g\cdot x)=g\,f(x)\,g^{\top}$ for all $g\in G$. This controlled setting allows precise evaluation of symmetry recovery.
6.2.2 Task 2: Moment of Inertia Prediction
Given $n$ point masses $m_{i}$ at positions $x_{i}\in\mathbb{R}^{3}$, we predict the inertia tensor:
$f(x,m)=\sum_{i=1}^{n}m_{i}\left(\|x_{i}\|^{2}I-x_{i}x_{i}^{\top}\right).$ (16)
This physically meaningful function is equivariant under the group $O(3)$:
$f(g\cdot x,m)=g\,f(x,m)\,g^{\top},\quad\forall g\in O(3).$ (17)
This task tests the ability to recover well-known physical symmetries from data.
6.2.3 Task 3: Top Quark Tagging
We classify jets as originating from top quarks versus light quarks using jet constituent four-vectors. The classification label is invariant under Lorentz transformations corresponding to the group
$G=O(1,3),\quad A=\eta=\mathrm{diag}(1,-1,-1,-1).$ (18)
This high-dimensional real-world task evaluates Lorentz invariance discovery in particle physics applications.
6.3 Evaluation Metrics
We assess performance using three complementary metrics:
Prediction Accuracy.
We measure the mean squared error (MSE) for regression tasks (Tasks 1–2) and the classification accuracy for the tagging task (Task 3). This quantifies how well models approximate the target function.
Quadratic Form Recovery.
We compute the cosine similarity between the learned and true quadratic forms:
$\cos(A_{0},A_{\text{learnt}})=\frac{\langle A_{\text{learnt}},A_{0}\rangle_{F}}{\|A_{\text{learnt}}\|_{F}\,\|A_{0}\|_{F}},$ (19)
where $\langle\cdot,\cdot\rangle_{F}$ denotes the Frobenius inner product. For $G$-Ortho-Nets, AA is learned directly; for LieGG and Augerino, we recover $A$ from the learned generators $X$ via the constraint $X^{\top}A+AX=0$.
Lie Algebra Comparison.
We compare the learned Lie algebra $\mathfrak{g}_{\text{learnt}}$ with the true Lie algebra $\mathfrak{g}_{0}$ using the principal angles. Given orthonormal bases for both subspaces, the principal angles $\theta_{1}\leq\theta_{2}\leq\cdots\leq\theta_{k}\in[0,\pi/2]$ characterize the geometric relationship between them. We compute the projection distance as:
$d_{\text{PA}}(\mathfrak{g}_{0},\mathfrak{g}_{\text{learnt}})=\sqrt{\sum_{i=1}^{k}\sin^{2}(\theta_{i})},$ (20)
which measures the overall misalignment between the two Lie algebra subspaces. A value of 0 indicates perfect alignment, while larger values indicate greater deviation.
6.4 Results
6.4.1 Quantitative results
Tables 3 and 4 summarize performance across all tasks. We evaluate prediction quality (MSE or accuracy), recovery of the quadratic form via $\cos(A_{0},A_{\text{learnt}})$, and Lie-algebra alignment using projection distance $d_{\text{PA}}(\mathfrak{g}_{0},\mathfrak{g}_{\text{learnt}})$ (principal angles). While LieGG achieves competitive performance in certain cases, it does not yield an interpretable prediction model and requires additional post-hoc analysis to extract Lie algebra generators.
6.4.2 Qualitative results
Figure 3 compares ground-truth and learned quadratic forms for the top-tagging and synthetic tasks, demonstrating faithful recovery of the target structure. $G$-Ortho-Nets achieve strong performance across both metrics and tasks. By explicitly parameterizing the quadratic form $A$, our method enables direct symmetry discovery, rather than inferring it indirectly from generators. On the top-tagging benchmark in particular, it attains state-of-the-art accuracy while most closely recovering the Lorentz structure, highlighting that encoding physically meaningful symmetries can improve generalization in high-energy physics.
Table 3: Regression tasks. Best results in bold. $\cos(A_{0},A_{\text{learnt}})$ is cosine similarity between true and learned AA; $d_{\text{PA}}(\mathfrak{g}_{0},\mathfrak{g}_{\text{learnt}})$ is projection distance from principal angles.
Task Method Pred. Error $\cos(A_{0},A_{\text{learnt}})$ $d_{\text{PA}}(\mathfrak{g}_{0},\mathfrak{g}_{\text{learnt}})$
Synthetic $G$-Ortho-Nets $1.80\times 10^{-3}$ 0.99 0.24
LieGG $\mathbf{7.53\times 10^{-4}}$ 0.93 1.80
Augerino $2.63\times 10^{0}$ 0.30 1.99
Inertia $G$-Ortho-Nets $\mathbf{1.88\times 10^{-3}}$ 0.99 0.15
LieGG $2.61\times 10^{-3}$ 0.99 0.06
Augerino $7.86\times 10^{-3}$ 0.57 1.26
Table 4: Top-tagging classification. Best results in bold.
Task Method Accuracy $\cos(A_{0},A_{\text{learnt}})$ $d_{\text{PA}}(\mathfrak{g}_{0},\mathfrak{g}_{\text{learnt}})$
Top tagging $G$-Ortho-Nets 90.44 0.9999 0.0214
LieGG 83.10 0.9991 1.0012
Augerino 51.80 0.5000 2.1830
(a) Ground-truth AA
(b) Learned AA
(c) Ground-truth AA
(d) Learned AA
Figure 3: Comparison of ground-truth and learned AA matrices for top-tagging (top) and synthetic regression (bottom) using $G$-Ortho-Nets.
Please refer to the appendix for additional experiments and analysis.
@@ 6.5 Limitations
Our framework is currently limited to orthogonal groups, as these are exactly the transformations that preserve quadratic forms. While this restriction excludes more general classes such as affine or diffeomorphic symmetries, it provides a clear advantage: interpretability. By explicitly learning the quadratic form $A$, our model yields a physically meaningful and directly interpretable predictor, rather than only an implicit or post-hoc symmetry characterization. This trade-off mirrors the broader tension in symmetry learning between generality and interpretability. We view our approach as complementary to methods such as LieGAN, Augerino or LIG, which target broader classes but do not produce usable equivariant predictors.
@@ 7 Conclusion
We presented a quadratic-form framework for learning group-equivariant functions, focusing on orthogonal groups. We proved a canonical decomposition into norm-invariant and scale-invariant components, and extended it to tuples under diagonal actions, where the equivariant map splits into an angular term (from the normalized first vector) and a scale-invariant term (from the full Gram matrix), thereby capturing interdependencies while preserving symmetry. Experiments across synthetic and physics-inspired tasks showed reliable symmetry discovery and effective equivariant learning. Future work includes extending beyond orthogonal groups and accommodating richer group actions.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes] We provide a clear mathematical formulation of the problem and describe the proposed model and assumptions
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [No] We do not provide a formal complexity analysis, but we discuss empirical efficiency.
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [No] Source code is not included in this submission, but we plan to release it in the camera-ready version.
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes] All assumptions underlying our theoretical results are explicitly stated
    2. (b)
      Complete proofs of all theoretical results. [Yes] Complete proofs are provided in the Appendix.
    3. (c)
      Clear explanations of any assumptions. [Yes] We provide clear explanations of all assumptions alongside the results.
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [No] We do include code, data to reproduce the main experimental results.
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes] Training details such as data splits and hyperparameters are described in the supplementary section.
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [No] We do not report evaluation metrics with error bars across multiple random seeds.
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes] The computing infrastructure, including GPU specifications are described in the supplementary section.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes] We cite all assets used in our work with proper references.
    2. (b)
      The license information of the assets, if applicable. [Not Applicable] License information of the used assets is included where applicable.
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Not Applicable] We do not release new assets in this work.
    4. (d)
      Information about consent from data providers/curators. [Not Applicable] No data requiring consent was collected.
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable] The work does not involve sensitive or personally identifiable content.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable] No human participants were involved.
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable] No IRB approval was required.
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable] No participant compensation was involved.
Supplementary Material
Index
  • $\triangleright$
    Additional Experiments ........................................................................................................................................................................A
  • $\triangleright$
    Notation and Standing Assumptions ........................................................................................................................................................................B
  • $\triangleright$
    Additional Results and Complete Proofs ........................................................................................................................................................................C
    • Canonical Alignment in the Positive Definite Case ........................................................................................................................................................................C.1
    • Canonical Alignment in the Positive Semi-Definite Case ........................................................................................................................................................................C.2
    • Canonical Alignment in the Indefinite Case ........................................................................................................................................................................C.3
    • Orbit Space and Quadratic-Form Norm Correspondence ........................................................................................................................................................................C.4
    • Decomposition into Scale- and Norm-Invariant Components ........................................................................................................................................................................C.5
    • Regularity conditions ........................................................................................................................................................................C.6
    • Extension to Diagonal action ........................................................................................................................................................................C.7
    • Householder reflection ........................................................................................................................................................................C.8
Appendix A Additional Experiments
A.1 Noise robustness
To evaluate the robustness of both methods to label corruption, we systematically vary the standard deviation of Gaussian noise added to the training labels. The underlying task is a synthetic polynomial regression problem, where the ground-truth function is given by $f(x)=9(xx^{\top}A)^{2}+2(xx^{\top}A)$. The goal is to assess how prediction accuracy degrades under increasing noise levels and to compare the noise tolerance of LieGG and our proposed $G$-Ortho-Nets. Table 5 reports the prediction error across different noise levels, highlighting the relative stability of each method.
Table 5: Comparison of $G$-Ortho-Nets and LieGG under varying label noise levels. Lower prediction error (MSE) and projection distance $d_{\text{PA}}$ indicate better performance, while higher cosine similarity $\cos(A_{0},A_{\text{learnt}})$ indicates closer alignment with the ground-truth quadratic form.
Std dev Method Pred. Error (MSE) $\boldsymbol{\cos(A_{0},A_{\text{learnt}})}$ $\boldsymbol{d_{\text{PA}}(\mathfrak{g}_{0},\mathfrak{g}_{\text{learnt}})}$
1.0 $G$-Ortho-Nets 1.0161 0.9778 0.3504
LieGG 43.7929 0.2900 1.7526
0.5 $G$-Ortho-Nets 0.2633 0.9882 0.2872
LieGG 10.9812 0.2468 1.7108
The results in Table 5 demonstrate the superior robustness and structural fidelity of the proposed $G$-Ortho-Nets compared to the LieGG baseline under increasing label noise. Even when subjected to substantial noise ($\sigma=1$), $G$-Ortho-Nets maintain a remarkably low prediction error (MSE $\approx 1.0$) and a high cosine similarity with the ground-truth quadratic form ($\cos(A_{0},A_{\text{learned}})\approx 0.98$), indicating that the learned symmetry remains closely aligned with the true underlying structure. In contrast, the LieGG method exhibits severe degradation, with prediction errors exceeding two orders of magnitude and a sharp drop in alignment metrics. These findings highlight that explicitly parameterizing the quadratic form and enforcing equivariance through the proposed decomposition enables stable learning of both the predictive function and the latent symmetry, even in the presence of substantial label corruption.
Appendix B Notation and Standing Assumptions
  • Quadratic form and group definition. Let $A\in\mathbb{R}^{n\times n}$ be a symmetric matrix. Define the generalized orthogonal group
    $G\;=\;\{\,Q\in\mathrm{GL}_{n}\;:\;Q^{\top}AQ=A\,\}.$
  • Indefinite norm. For $x\in\mathbb{R}^{n}$, define the $A$-norm (or pseudo-norm)
    $\|x\|_{A}\;=\;\operatorname{sign}(x^{\top}Ax)\,\sqrt{|x^{\top}Ax|}.$
    This norm generalizes the Euclidean case and may take positive or negative values depending on the signature of $A$.
  • Null cone. We exclude the null cone
    $U\;=\;\{\,x\in\mathbb{R}^{n}:x^{\top}Ax=0\,\},$
    which has Lebesgue measure zero. All maps and actions below are defined on $\mathbb{R}^{n}\setminus U$.
  • Group action on inputs. The group $G$ acts on the input space by the standard linear action
    $Q\cdot x\;=\;Qx.$
  • Group action on outputs. When the output space is matrix-valued, we consider the conjugation action
    $Q\cdot Y\;=\;QYQ^{-1},$
    which preserves the natural geometric structure of the output space.
Appendix C Additional Results and Complete Proofs